a woman backs her car out of her garage with an acceleration of 2.80m/s^2. how long does it take her to reach a speed of 4.00m/s?
1
Expert's answer
2020-10-01T10:10:35-0400

We can find the time that woman takes to reach a speed of 4.00 m/s from the kinematic equation:


"v = v_0 + at,"

here, "v_0 = 0 \ \dfrac{m}{s}" is the initial velocity of the car, "v = 4.0 \ \dfrac{m}{s}" is the final velocity of the car, "a = 2.80 \ \dfrac{m}{s^2}" is the acceleration of the car and "t" is the time that woman takes to reach a speed of 4.00 m/s.

Then, we get:


"t = \dfrac{v-v_0}{a} = \dfrac{4.0 \ \dfrac{m}{s}-0 \ \dfrac{m}{s}}{2.80 \ \dfrac{m}{s^2}} = 1.43 \ s."

Answer:

"t = 1.43 \ s."